top of page

DEPOT REPORT WEDNESDAY 19th JUNE.

VLO. New working member Graham Isaac attended for his induction and stayed for the rest of the day helping out with the promise to return soon.


55009.

No 9 was returned to its own identity today, with The Black Watch plates and numbers removed from one side, rest to follow on Saturday, picture Alex Stojanovic. The 'A' side main air tank was removed to be loaned to 19 to help with testing next week.


55015.

The 'Delectricians' made a start today on checking all the electrical circuits. A few gremlins were found but nothing too serious, work to continue on Saturday. A start was made on a very light flat rub down before a first coat of varnish, A side fuel tanks were given their first coat, see picture attached from Alex Stojanovic.


55019.

As mentioned above a main air tank was borrowed from 9 and fitted to 19 today in advance of one of the 2 sent away being collected on Thursday and to be fitted Saturday to allow testing to proceed as the remaining tank is not due back until late July.

We also had a fabricator on site to replace both nose end footsteps with newly, offsite fabricated ones, picture Mike Hallam-Rudd.


Power Units.

Work continued on 430 with work centred on the hoses inside the 'triangle'.





404 views0 comments

Recent Posts

See All

Opmerkingen


bottom of page